Solve by using the substitution method.
y=-3x-1
2x-3y=-8

Answer by HyperBrain(694)

We need to substitute y=-3x-1 to 2x-3y=-8.
2x-3(-3x-1)=-8
2x-(-9x-3)=-8
2x+9x+3=-8
11x+3=-8
11x=-11
x=-1
But, y=-3x-1. So,
y=-3(-1)-1=3-1=2
Therefore, x=-1 and y=2
Power up,
HyperBrain!
pastime blogtimehere
use less plasticnow